math

Find a polynomial function of degree 3 with real coefficients that satisfies the given conditions.

Zero of - 4 having multiplicity 3; f(1) = 25.
Answer #1

Here is the detailed answer for the question.

Given that zero to -4 having multiplicity 3

So, f(x)=a(x+4)^3

we know that f(1)=25

placing x value as 1

f(1)=a(1+4)^3

25=125a

a=1/5

So, the polynomial with degree 3 is f(x)=((x+4)^3)/5

answered by: anonymous